Could someone please help me? I am having trouble with this homework question.
A silversmith mixed 30 g of a 60% silver alloy with 70 g of a 20% silver alloy. What is the percent concentration of the resulting alloy? Thanks for any help. I'm not sure how to even set it up right.

Answer by ankor@dixie-net.com(22740)

A silversmith mixed 30 g of a 60% silver alloy with 70 g of a 20% silver alloy. What is the percent concentration of the resulting alloy?
:
This is typical mixture problem, learn this method and you can do most of them.
:
We know that amt of the resulting alloy will be the total of the two inputs
30 + 70 = 100 gram
:
Convert % to decimals
:
Let x = per cent (in decimals) of the resulting alloy:
:
.60(30) + .20(70) = x(100)
18 + 14 = 100x
32 = 100x
x = 32/100 = .32 = 32%
:
Pretty easy, right?